OpenStudy (anonymous):

since the sides of a square are congruent write an equation that will help you find the length of a

OpenStudy (anonymous):



OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

all sides are equal in a square

ganeshie8 (ganeshie8):

if one side is \(a\), all other sides also will be \(a\)

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

use pythagora

OpenStudy (anonymous):

So would the equation be a^2+a^2=55^2

ganeshie8 (ganeshie8):

perfect ! thats the equation, good work !

OpenStudy (anonymous):

ohh i cant believe i didn't get that

OpenStudy (anonymous):

the next question tells me to solve the equation and round to the nearest 10th

OpenStudy (anonymous):

a^2+a^2=55^2 a^2+a^2=3025

ganeshie8 (ganeshie8):

a^2+a^2=55^2 2a^2 = 55^2 a^2 = 55^2/2 a = 55/sqrt(2)

ganeshie8 (ganeshie8):

use ur calculator

OpenStudy (anonymous):

would i get 38.89

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

then i have to round to the nearest 10th which would be

OpenStudy (anonymous):

38.90

ganeshie8 (ganeshie8):

38.9

OpenStudy (anonymous):

yeah i just realized that

ganeshie8 (ganeshie8):

:)